OpenStudy (anonymous):

How would you do a partial fraction decomposition for (3x+4)/(x^3-2x-4). Anybody?

OpenStudy (anonymous):

\[\frac{3 x+4}{x^3-2 x-4}=\frac{1}{x-2}-\frac{x+1}{x^2+2 x+2} \]